OpenStudy (anonymous):

Q=I^2Rt/J Solve for R

OpenStudy (lynfran):

\[Q=\frac{ I^2Rt }{ J }\]like this??

OpenStudy (anonymous):

Yes

OpenStudy (lynfran):

first step.. multiply both sides by J''

OpenStudy (lynfran):

what do u get?

OpenStudy (anonymous):

J cancels out

OpenStudy (anonymous):

So you are left with the rest

OpenStudy (lynfran):

yes on the right side of the equation and we r left with \[QJ=I^2Rt\]

OpenStudy (lynfran):

next step divide both side by ....do u know?

OpenStudy (anonymous):

nope

OpenStudy (lynfran):

we need the R term alone so we can divide both sides by I^2t

OpenStudy (lynfran):

\[\frac{ QJ }{ I^2t }=\frac{ I^2Rt }{ I^2t }\]

OpenStudy (lynfran):

so the I^2t cancel out leaving R

OpenStudy (lynfran):

Do u see that?

OpenStudy (anonymous):

Ah yes yes I see

OpenStudy (lynfran):

ok

OpenStudy (anonymous):

So, what do you do after that?

OpenStudy (lynfran):

we finish

OpenStudy (lynfran):